Manny is 9 years older than Charlie. Four years ago the sum of their ages was 17 years. In how many years will Charlie be 18 years of age? Please help me
Answer by macston(5194)

.
C=Charlie's age; M=Manny's age=C+9
.
(C-4)+(M-4)=17
C-4+(C+9)-4=17
2C+1=17
2C=16
C=16 Charlie is 16 years old.
18yrs-16yrs=2yrs
.
ANSWER: Charlie will be 18 years old in 2 years.